# Gemma Scope: Open Sparse Autoencoders Everywhere All At Once on Gemma 2

### v1

@@ 1 Introduction
These are several lines of evidence that suggest that a significant fraction of the internal activations of language models are sparse, linear combination of vectors, each corresponding to meaningful features (Elhage et al., 2022; Gurnee et al., 2023; Olah et al., 2020; Park et al., 2023; Nanda et al., 2023a; Mikolov et al., 2013). But by default, it is difficult to identify which vectors are meaningful, or which meaningful vectors are present. Sparse autoencoders are a promising unsupervised approach to do this, and have been shown to often find causally relevant, interpretable directions (Bricken et al., 2023; Cunningham et al., 2023; Templeton et al., 2024; Gao et al., 2024; Marks et al., 2024). If this approach succeeds it could help unlock many of the hoped for applications of interpretability (Nanda, 2022; Olah, 2021; Hubinger, 2022), such as detecting and fixing hallucinations, being able to reliably explain and debug unexpected model behaviour and preventing deception or manipulation from autonomous AI agents.
@@ 2.2 JumpReLU SAEs
where $\lambda$ is the sparsity penalty coefficient. Since the L0 penalty and JumpReLU activation function are piecewise constant with respect to threshold parameters ${\boldsymbol{\theta}}$, we use straight-through estimators (STEs) to train ${\boldsymbol{\theta}}$, using the approach described in Rajamanoharan et al. (2024b). This introduces an additional hyperparameter, the kernel density estimator bandwidth $\varepsilon$, which controls the quality of the gradient estimates use to train the threshold parameters ${\boldsymbol{\theta}}$ [^5] A large value of $\varepsilon$ results in biased but low variance estimates, leading to SAEs with good sparsity but sub-optimal fidelity, whereas a low value of $\varepsilon$ results in high variance estimates that cause the threshold to fail to train at all, resulting in SAEs that fail to be sparse. We find through hyperparameter sweeps across multiple layers and sites that $\varepsilon=0.001$ provides a good trade-off (when SAE inputs are normalized to have a unit mean squared norm) and use this to train the SAEs released as part of Gemma Scope.
@@ 3.3.2 Data Pipeline
We store all collected activations on hard drives as raw bytes in shards of 10-20GiB. We use 32-bit precision in all our experiments. This means that storing 8B worth of activations for a single site and layer takes about 100TiB for Gemma 2 9B, or about 17PiB for all sites and layers of both Gemma 2 2B and 9B. The total amount is somewhat higher still, as we train some SAEs for 16B tokens and also train some SAEs on Gemma 2 27B, as well as having a generous buffer of additional tokens. While this is a significant amount of disk space, it is still cheaper than regenerating the data every time one wishes to train an SAE on it.
@@ 4.7 Impact of low precision inference
We train all SAEs in 32-bit floating point precision. It is common to make model inference less memory and compute intensive by reducing the precision at inference time. This is particularly important for applications like circuit analysis, where users may wish to splice several SAEs into a language model simultaneously. Fig. 10 comparse fidelity-versus-sparsity curves computed using float32 SAE and LM weights versus the same curves computed using bfloat16 SAE and LM weights, suggesting there is negligible impact in switching to bfloat16 for inference.

### v2

@@ 1 Introduction
There are several lines of evidence that suggest that a significant fraction of the internal activations of language models are sparse, linear combination of vectors, each corresponding to meaningful features (Elhage et al., 2022; Gurnee et al., 2023; Olah et al., 2020; Park et al., 2023; Nanda et al., 2023a; Mikolov et al., 2013). But by default, it is difficult to identify which vectors are meaningful, or which meaningful vectors are present. Sparse autoencoders are a promising unsupervised approach to do this, and have been shown to often find causally relevant, interpretable directions (Bricken et al., 2023; Cunningham et al., 2023; Templeton et al., 2024; Gao et al., 2024; Marks et al., 2024). If this approach succeeds it could help unlock many of the hoped for applications of interpretability (Nanda, 2022; Olah, 2021; Hubinger, 2022), such as detecting and fixing hallucinations, being able to reliably explain and debug unexpected model behaviour and preventing deception or manipulation from autonomous AI agents.
@@ 2.2 JumpReLU SAEs
where $\lambda$ is the sparsity penalty coefficient. Since the L0 penalty and JumpReLU activation function are piecewise constant with respect to threshold parameters ${\boldsymbol{\theta}}$, we use straight-through estimators (STEs) to train ${\boldsymbol{\theta}}$, using the approach described in Rajamanoharan et al. (2024b). This introduces an additional hyperparameter, the kernel density estimator bandwidth $\varepsilon$, which controls the quality of the gradient estimates used to train the threshold parameters ${\boldsymbol{\theta}}$ [^5] A large value of $\varepsilon$ results in biased but low variance estimates, leading to SAEs with good sparsity but sub-optimal fidelity, whereas a low value of $\varepsilon$ results in high variance estimates that cause the threshold to fail to train at all, resulting in SAEs that fail to be sparse. We find through hyperparameter sweeps across multiple layers and sites that $\varepsilon=0.001$ provides a good trade-off (when SAE inputs are normalized to have a unit mean squared norm) and use this to train the SAEs released as part of Gemma Scope.
@@ 3.3.2 Data Pipeline
We store all collected activations on hard drives as raw bytes in shards of 10-20GiB. We use 32-bit precision in all our experiments. This means that storing 8B worth of activations for a single site and layer takes about 100TiB for Gemma 2 9B, or about 17PiB for all sites and layers of both Gemma 2 2B and 9B. The total amount is somewhat higher still, as we train some SAEs for 16B tokens and also train some SAEs on Gemma 2 27B, as well as having a generous buffer of additional tokens. While this is a significant amount of disk space, it is still cheaper than regenerating the data every time one wishes to train an SAE on it. Concretely, in our calculations we find that storing activations for 10-100 days is typically at least an order of magnitude cheaper than regenerating them one additional time. The exact numbers depend on the model used and the specifics of the infrastructure, but we expect this relationship to hold true in general. If there is a hard limit on the amount of disk space available, however, or if fast disk I/O can not be provided (see next paragraph), then this will favor on-the-fly generation instead. This would also be the case if the exact hyperparameter combinations were known in advance. In practice, we find it advantageous for research iteration speed to be able to sweep sparsity independently from other hyperparameters and to retrain SAEs at will.
@@ 4.7 Impact of low precision inference
We train all SAEs in 32-bit floating point precision. It is common to make model inference less memory and compute intensive by reducing the precision at inference time. This is particularly important for applications like circuit analysis, where users may wish to splice several SAEs into a language model simultaneously. Fig. 10 compares fidelity-versus-sparsity curves computed using float32 SAE and LM weights versus the same curves computed using bfloat16 SAE and LM weights, suggesting there is negligible impact in switching to bfloat16 for inference.